# G2D2: Gradient-guided Discrete Diffusion for image inverse problem solving

### v1

@@ 3 Gradient-guided discrete diffusion, G2D2
Besides the lack of a straightforward mechanism to propagate gradients of the loss function through the generative process, a preliminary study reveals another main issue of directly applying the graphical model of a general mask-absorbing discrete diffusion model to sampling in an inverse-problem context. Figure 2 shows images decoded from $\mathbf{z}_{0}$, which are sampled from the denoising model $\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})$ conditioned on the intermediate noisy discrete latent $\mathbf{z}_{90}$ or $\mathbf{z}_{80}$, along with the generated image from the full reverse process. These images are generated using a pre-trained VQ-Diffusion model (Gu et al., 2022) with the prompt “A face of monkey” (where $T=100$). The results indicate that the majority of the image structure is determined within the initial approximately 10% of the steps.
Figure 2: At time step $t$, $\mathbf{z}_{0}$ is sampled from the prior model $\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})$ and decoded. These results are generated with prompt “A face of monkey”, without any guidance. By initial $\sim$10 steps ($t=90$), coarse structure of image has already been determined.
This observation highlights a problem for commonly used mask-absorbing discrete diffusion models in the context of inverse problems. The issue arises from the definition of $Q_{t}$ in (2), where $\beta_{t}$ is usually set to be extremely small. In the forward process, therefore, unmasked tokens are highly likely to either remain as identical unmasked tokens or transition to masked tokens. Masked tokens also remain unchanged thereafter.
This characteristic indicates that in the reverse process, the probability of unmasked tokens reverting to masked tokens or transitioning to different unmasked tokens is negligible. Consequently, when sampling to satisfy the measurement model, errors occurring in the early stages of sampling become nearly impossible to correct in the later phases.
One solution to this problem is the “re-masking” operation, which reverts unmasked tokens back to masked tokens. Similar approaches have been used with discrete predictor-corrector methods (Lezama et al., 2023) and predictor-corrector techniques for continuous-time discrete diffusion (Campbell et al., 2022; Zhao et al., 2024) to improve image-generation quality. However, those that involve reversing time steps can increase computational complexity. To address this, we demonstrate that by considering a noise process that is independent at each time step and different from the one used during the training of the prior, we can naturally resolve the inherent issues associated with mask-absorbing discrete diffusion models. We also show that despite this difference in the noise process, the prior model can still be used within our framework.
3.1 Star-shaped noise process for inverse problem solving
Figure 3: Graphical models using Markov noise process (upper) and the star-shaped noise process (lower)
Inspired by Okhotin et al. (2024) and Zhang et al. (2024), G2D2 employs the star-shaped noise process. Figure 3 illustrates the differences between the Markov forward noise process (upper), which is used in general discrete diffusion models, and the star-shaped noise process (lower), both incorporating the relationship with the measurement $\mathbf{y}$. In the star-shaped noise process, the noisy discrete latents $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}$ are conditionally independent given $\mathbf{z}_{0}$. We assume that the distribution $q(\mathbf{z}_{t}|\mathbf{z}_{0})$ adopts the same form as the original forward Markov process, specifically $q(z_{t,i}|\mathbf{z}_{0})=\bm{v}^{\mathsf{T}}(z_{t,i})\overline{Q}_{t}\bm{v}(z_{0,i})$, with $\overline{Q}_{t}=Q_{t}\cdots Q_{1}$.
We aim to sample from the posterior $q(\mathbf{z}_{0}|\mathbf{y})$ given the measurement $\mathbf{y}$ based on this graphical model. Given that the transformation from $\mathbf{z}_{0}$ to $\mathbf{x}_{0}$ is nearly deterministic in general decoders, we omit the random variable $\mathbf{x}_{0}$ in the subsequent discussion for simplicity.
To discuss the implementation of the sampling method based on the graphical model, we first introduce the conditional joint distribution $q_{\text{sampling}}(\mathbf{z}_{0:T}|\mathbf{y})=q(\mathbf{z}_{T}|\mathbf{y})\prod_{t=1}^{T}q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$. This conditional joint distribution has several properties:
@@ 3.1 Star-shaped noise process for inverse problem solving
1. A single step of $q_{\text{sampling}}$ inherently enables the “re-masking” operation.
@@
In the star-shaped noise process, the positions of mask tokens in $\mathbf{z}_{t-1}$ and $\mathbf{z}_{t}$ are mutually independent and uncorrelated. Consequently, the conditional distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$ enables a “re-masking” operation, wherein unmasked tokens present in $\mathbf{z}_{t}$ can become masked tokens in $\mathbf{z}_{t}$. This property suggests that in mask-absorbing discrete diffusion, errors that occur in the initial stages of sampling can be corrected in subsequent steps, which provides an advantage when solving inverse problems.
2. The marginal distribution $q_{\text{sampling}}(\mathbf{z}_{0}|\mathbf{y})$ is identical to the target distribution $q(\mathbf{z}_{0}|\mathbf{y})$ if the conditional distributions $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$ are correctly specified based on the graphical model of the star-shaped noise process.
The statement and proof are provided in the Appendix. This indicates that if sampling from each step $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$ in $q_{\text{sampling}}$ is feasible, then it is possible to sample from the target marginal posterior $q(\mathbf{z}_{0}|\mathbf{y})$.
@@
3. The conditional joint distribution $q_{\text{sampling}}$ differs from that of the graphical model, i.e., $q_{\text{sampling}}(\mathbf{z}_{0:T}|\mathbf{y})\neq q(\mathbf{z}_{0:T}|\mathbf{y})$.
@@
In a standard Markov forward process, the decomposition of the joint distribution in a star-shaped noise process takes the form $q(\mathbf{z}_{0:T}|\mathbf{y})=q(\mathbf{z}_{T}|\mathbf{y})\prod_{t=1}^{T}q(\mathbf{z}_{t-1}|\mathbf{z}_{t:T},\mathbf{y})$. However, $q_{\text{sampling}}$ deviates from this formulation by disregarding the dependencies on larger time steps, $\mathbf{z}_{t+1:T}$.
@@
In subsequent sections, we introduce a variational distribution to approximate $q_{\text{sampling}}$, which inherently enables the re-masking operation based on Property 1. As established by Property 3, the joint distribution of $q_{\text{sampling}}$ differs from that of the star-shaped noise process graphical model. Nevertheless, Property 2 ensures that they share identical marginal distributions. Moreover, given that $q_{\text{sampling}}$ focuses on only two adjacent variables, we can formulate an algorithm to approximate its distribution using a variational approach.
3.2 G2D2 based on star-shaped noise process
@@
Based on the discussion in the previous section, we aim to implement $q_{\text{sampling}}$ on the graphical model of the star-shaped noise process, which inherently incorporates a re-masking process. Specifically, we introduce a variational distribution $p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})$ to approximate $q_{\text{sampling}}(\mathbf{z}_{0:T}|\mathbf{y})$, with the ultimate goal of ensuring that the marginal distribution $p_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{y})$ approximates the true posterior $q(\mathbf{z}_{0}|\mathbf{y})$. The distribution $p_{\bm{\alpha}}$ is decomposed as
$p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})=\sum_{\mathbf{z}_{0}}q(\mathbf{z}_{t-1}|\mathbf{z}_{0})\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y}),$ (5)
@@
where $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ is a categorical distribution parameterized by $\bm{\alpha}\in\mathbb{R}^{T\times d_{\mathbf{z}}\times K}$, defined as $\tilde{p}_{\bm{\alpha}}(z_{0,i}|\mathbf{z}_{t},\mathbf{y})=\text{Cat}\left(z_{0,i};\bm{\alpha}_{t,i,\cdot}\right)$, i.e., $\tilde{p}_{\bm{\alpha}}(z_{0,i}=k|\mathbf{z}_{t},\mathbf{y})=\alpha_{t,i,k}$. This decomposition stems from the fact that the distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$ can be expressed as $\sum_{\mathbf{z}_{0}}q(\mathbf{z}_{t-1}|\mathbf{z}_{0})q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ based on the conditional independence. Note that both $q(\mathbf{z}_{t-1}|\mathbf{z}_{0})$ and $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ have a mean field structure with independent categorical distributions across dimensions. Consequently, $p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$, obtained by marginalizing over $\mathbf{z}_{0}$, inherits this mean field property. For notational convenience, we denote the slice of distribution parameter $\bm{\alpha}$ at time step $t$ as $\bm{\alpha}_{t}\in\mathbb{R}^{d_{\mathbf{z}}\times K}$.
@@
In G2D2, the variational distribution $p_{\bm{\alpha}}$ is obtained by optimizing an objective function derived from the following theorem:
Theorem 3.1.
@@
Let $p_{\bm{\alpha}}$ be a distribution with the parameterization given by the decomposition in (5). Then, for any measurement $\mathbf{y}$, the following inequality holds for the Kullback-Leibler (KL) divergence between the marginal distributions:
@@
$D_{\mathrm{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{y})\|q(\mathbf{z}_{0}|\mathbf{y})\right)\leq\sum_{t=1}^{T}\mathbb{E}_{\mathbf{z}_{t}\sim p_{\bm{\alpha}}(\mathbf{z}_{t}|\mathbf{y})}\left[D_{\mathrm{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)\right].$ (6)
The full definitions of the terms are provided in the Appendix.
@@
The proof is provided in the Appendix. Based on this inequality, we aim to minimize each term in the sum on the right-hand side. Since $\tilde{p}_{\bm{\alpha}}$ is a different categorical distribution at each $t$, we minimize $\bm{\alpha}$ for each time step, ultimately aiming to minimize the left-hand side. Each term on the right-hand side of (6) takes the following form:
Lemma 3.2.
@@
The KL divergence between the variational distribution $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ and true conditional $q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ can be decomposed into two terms:
@@
$D_{\mathrm{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)=D_{\mathrm{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q(\mathbf{z}_{0}|\mathbf{z}_{t})\right)-\mathbb{E}_{\mathbf{z}_{0}\sim\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}\left[\log q(\mathbf{y}|\mathbf{z}_{0})\right],$ (7)
The full definitions of these terms are provided in the Appendix.
This decomposition enables us to separately consider the fit to the prior and the consistency with the measurement data.
@@
The first term on the right-hand side of (7) remains intractable. However, it is important to note that the star-shaped noise process shares the conditional distribution $q(\mathbf{z}_{t}|\mathbf{z}_{0})$ with the original Markov noise process. Consequently, the reverse conditional distribution $q(\mathbf{z}_{0}|\mathbf{z}_{t})$ will also be identical for both processes. Since the prior of the pre-trained discrete diffusion models is trained to approximate this distribution, we substitute this prior model $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t})$ for $q(\mathbf{z}_{0}|\mathbf{z}_{t})$ into the objective function of (7). This substitution transforms the term into a KL divergence between two categorical distributions, enabling the computation of gradients with respect to the parameter $\bm{\alpha}$.
@@
Based on Theorem 3.1 and Lemma 3.2, G2D2 optimizes the parameter $\bm{\alpha}$ of $p_{\bm{\alpha}}$ for $t=T,\ldots,1$ while sequentially sampling $\mathbf{z}_{0:T}$. In the optimization step, any continuous optimization method, such as Adam (Kingma, 2014), can be used. Implementation considerations are discussed in the following section. This algorithm is detailed in Algorithm 1, and G2D2 is illustrated in Figure 1.
@@
5:         Initialize: $\bm{\alpha}_{t}=\log\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})$

### v2

@@ 3 Gradient-Guided Discrete Diffusion, G2D2
Figure 2: Empirical demonstration of mask-absorbing characteristics in discrete diffusion models. All images are generated using the text prompt “a face of monkey”. Top row: Diverse samples from text-conditioned sampling ($p_{\theta}(\mathbf{z}_{0})$). Middle row: Multiple samples conditioned on the same $\mathbf{z}_{t=90}$ ($p_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t=90})$). Bottom row: Multiple samples conditioned on the same $\mathbf{z}_{t=80}$ ($p_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t=80})$). This demonstrates that in mask-absorbing processes, image structure is largely determined in early sampling steps, with decreasing diversity as sampling progresses, making it difficult to correct inconsistencies with measurement data during inverse problem solving.
In this section, we propose Gradient-Guided Discrete Diffusion (G2D2), an inverse problem solving method that leverages discrete diffusion models as priors. Our approach addresses two key challenges: the non-differentiability of discrete diffusion models and the limitations imposed by widely-used mask-absorbing noise processes. First, we introduce a star-shaped noise process to address the latter issue and define a more tractable variant of the star-shaped noise process distribution, which we call the star-decomposed distribution. We then address the former challenge of non-differentiability by bridging the gap between discrete and continuous domains through optimizing a variational distribution with continuous relaxation. These methodological components are integrated into a unified, practical inference algorithm.
3.1 Star-shaped noise process for enabling inherent re-masking
A key challenge in applying discrete diffusion models to inverse problems is addressing the limitations of mask-absorbing noise processes. As described in Section 2.1, discrete diffusion models commonly adopt a mask-absorbing process, where in the forward process, an unmasked token either remains the same or transitions to a mask token. While this design leads to higher performance in standard generation tasks, it poses a significant constraint in solving inverse problems. Specifically, in the reverse process (i.e., the generative process), once an unmasked token has been set at an early stage of sampling, the probability of it reverting to a masked state or changing to a different token becomes extremely low (see Figure 3(b)), making it difficult to correct errors made in the early stages.
Figure 3: (a) Graphical model comparison between Markov noise process (top) and star-shaped noise process (bottom). The star-shaped model introduces conditional independence between $\mathbf{z}_{t}$ variables given $\mathbf{z}_{0}$, enabling more flexible error correction during sampling. (b) Comparison between mask-absorbing Markov process (left) and mask-absorbing star-shaped process (right). In the Markov process, unmasked tokens cannot revert to [MASK], while the star-shaped process allows tokens to naturally revert to [MASK] state (inherent “re-masking”), facilitating error correction in later sampling steps.
Figure 2 demonstrates this characteristic: when generating images using discrete diffusion models that employ mask-absorbing process, most of the image structure is determined in the initial stages, suggesting that in inverse problem scenarios, there is little opportunity to correct inconsistencies with the measurement $\mathbf{y}$ during sampling. One solution to this problem is the “re-masking” operation, which reverts unmasked tokens to masked tokens. This approach has been implemented in discrete predictor-corrector methods (Lezama et al., 2023) and predictor-corrector techniques for continuous-time discrete diffusion (Campbell et al., 2022; Zhao et al., 2024) to improve image generation quality. However, these methods often involve reverting the sampler to earlier time steps and can be computationally expensive.
Here, drawing inspiration from Okhotin et al. (2024) and Zhang et al. (2024), we adopt a star-shaped noise process that decouples each time step $t$ from previous steps, which addresses the aforementioned issue of early-stage errors becoming permanently “frozen.” In this star-shaped noise process, the noisy variables $\mathbf{z}_{1},\dots,\mathbf{z}_{T}$ are conditionally independent given $\mathbf{z}_{0}$. Formally, we assume $q_{\text{star}}(\mathbf{z}_{1:T}|\mathbf{z}_{0})=\prod_{t=1}^{T}q_{\text{star}}(\mathbf{z}_{t}|\mathbf{z}_{0})=\prod_{t=1}^{T}\prod_{i=1}^{d_{\mathbf{z}}}\Bigl[\bm{v}^{\mathsf{T}}(z_{t,i})\,\overline{Q}_{t}\,\bm{v}(z_{0,i})\Bigr]$, where the matrix $\overline{Q}_{t}=Q_{t}\cdots Q_{1}$ is the cumulative transition from the original Markov forward process in discrete diffusion, but the forward graphical model is star-shaped, as shown in Figure 3(a).
Compared to the original Markov forward process, the star-shaped noise process maintains the same conditional marginal distribution $q_{\text{star}}(\mathbf{z}_{t}|\mathbf{z}_{0})$, yet the locations of [MASK] tokens are uncorrelated between $\mathbf{z}_{t}$ corresponding to different time steps. In the Markov noise process, once a token is [MASK]ed, it remains so, and in subsequent steps, only unmasked tokens can transition to [MASK] tokens. In contrast, in the star-shaped noise process, each noisy token $\mathbf{z}_{t}$ is viewed as being generated directly from $\mathbf{z}_{0}$ rather than from $\mathbf{z}_{t-1}$, making the [MASK] token positions independent between adjacent $\mathbf{z}_{t-1}$ and $\mathbf{z}_{t}$.
A key characteristic of the star-shaped process is observed in the reverse step $q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t})$. Unlike the Markov noise process with mask-absorbing state, which requires that tokens already unmasked remain fixed and prohibits their reversion to a masked state, the star-shaped formulation permits tokens to transition from an unmasked state back to a masked state within a single reverse step. This inherent “re-masking” operation enables flexible error correction during sampling, as illustrated in Figure 3(b), where tokens can naturally revert to the masked state without the need to roll back the sampler.
Our goal here, similar to other diffusion model-based inverse problem methods, is to sample from $q_{\text{star}}(\mathbf{z}_{0}|\mathbf{y})$ based on this graphical model. More specifically, we aim to develop a feasible algorithm for sampling from $q_{\text{star}}(\mathbf{z}_{0:T}|\mathbf{y})$. However, attempting to perform ancestral sampling from this distribution, such as by proceeding from $\mathbf{z}_{T}\rightarrow\mathbf{z}_{T-1}\rightarrow\cdots\rightarrow\mathbf{z}_{0}$, results in a posterior of the form $q_{\text{star}}(\mathbf{z}_{t}|\mathbf{z}_{t+1:T},\mathbf{y})$, which depends on all subsequent samples and is therefore not computationally feasible.
To address this challenge, we first introduce a proxy joint distribution, $q_{\text{star-decomp}}(\mathbf{z}_{0:T}|\mathbf{y})$, obtained by factorizing the star-shaped posterior into a product of local conditionals. Importantly, this preserves the marginal $q_{\text{star-decomp}}(\mathbf{z}_{0}|\mathbf{y})=q_{\text{star}}(\mathbf{z}_{0}|\mathbf{y})$ as we will discuss in the next subsection, so it can serve as a valid surrogate for $q_{\text{star}}$ while reducing the complexity of the graph structure so that dependencies exist only between consecutive time steps. Although the full joint $q_{\text{star-decomp}}$ is still intractable, this decomposition is a necessary preparatory step that leads to the variational approximation described in Sec. 3.3. The resulting algorithm separates the complex inter-dependencies between time steps while maintaining the essential properties of the star-shaped model, making it possible to efficiently sample from $q_{\text{star}}(\mathbf{z}_{0}|\mathbf{y})$.
3.2 Defining the decomposed star-shaped distribution $q_{\text{star-decomp}}$ which is a tractable variant of $q_{\text{star}}$
We define the star-decomposed distribution as follows:
$q_{\text{star-decomp}}(\mathbf{z}_{0:T}|\mathbf{y})=q_{\text{star}}(\mathbf{z}_{T}|\mathbf{y})\prod_{t=1}^{T}q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y}),$ (5)
where $q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$ represents the conditional distribution of $q_{\text{star}}$. This decomposition of the joint distribution resembles the Markovian reverse process of diffusion models, but it has the following properties:
@@ 3.1 Star-shaped noise process for enabling inherent re-masking
1. A single step of $q_{\text{star-decomp}}$ inherently enables the “re-masking” operation.
@@
In the star-shaped noise process, the positions of mask tokens in $\mathbf{z}_{t-1}$ and $\mathbf{z}_{t}$ are mutually independent and uncorrelated as shown in Figure 3(b). Consequently, the conditional distribution $q_{\text{star-decomp}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$ ($q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$) enables a “re-masking” operation, wherein unmasked tokens present in $\mathbf{z}_{t}$ can become masked tokens in $\mathbf{z}_{t-1}$. This property suggests that in mask-absorbing discrete diffusion, errors that occur in the initial stages of sampling can be corrected in subsequent steps, which provides an advantage when solving inverse problems.
2. The marginal distribution $q_{\text{star-decomp}}(\mathbf{z}_{0}|\mathbf{y})$ is identical to the target distribution $q_{\text{star}}(\mathbf{z}_{0}|\mathbf{y})$.
The statement and proof are provided in the Appendix. This suggests that for solving inverse problems, we do not necessarily need to sample from the joint distribution of $q_{\text{star}}$, but can instead aim to sample from the more tractable distribution $q_{\text{star-decomp}}$. Ultimately, we perform approximate sampling from the posterior by approximating this $q_{\text{star-decomp}}$ using a variational distribution.
@@
3. The conditional joint distribution of $q_{\text{star-decomp}}$ differs from that of $q_{\text{star}}$, i.e., $q_{\text{star-decomp}}(\mathbf{z}_{0:T}|\mathbf{y})\neq q_{\text{star}}(\mathbf{z}_{0:T}|\mathbf{y})$.
@@
The decomposition of the joint distribution of a star-shaped noise process takes the form $q_{\text{star}}(\mathbf{z}_{0:T}|\mathbf{y})=q_{\text{star}}(\mathbf{z}_{T}|\mathbf{y})\prod_{t=1}^{T}q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t:T},\mathbf{y})$. However, $q_{\text{star-decomp}}$ deviates from this formulation by disregarding the dependencies on larger time steps, $\mathbf{z}_{t+1:T}$.
@@
In subsequent sections, we introduce a variational distribution to approximate $q_{\text{star-decomp}}$, which possesses Property 1 that inherently enables the re-masking operation. As established by Property 3, the joint distribution of $q_{\text{star-decomp}}$ differs from that of the star-shaped noise process graphical model. Nevertheless, Property 2 guarantees that they share identical marginal distributions given the measurement $\mathbf{y}$, providing justification for our approach of targeting the more tractable $q_{\text{star-decomp}}$ for sampling. Furthermore, since $q_{\text{star-decomp}}$ focuses only on two adjacent variables, we can formulate a simple algorithm to approximate its distribution using a variational approach.
3.3 Variational approximation for feasible inference algorithm
@@
Based on the discussion in the previous section, we aim to implement $q_{\text{star-decomp}}$, which inherently incorporates a re-masking process for efficient inverse problem solving. However, since $q_{\text{star-decomp}}$ is still not tractable, we introduce a variational distribution $p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})=q_{\textnormal{star}}(\mathbf{z}_{T}|\mathbf{y})\prod_{t=1}^{T}p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$ to approximate $q_{\text{star-decomp}}(\mathbf{z}_{0:T}|\mathbf{y})$, with the ultimate goal of ensuring that the marginal distribution $p_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{y})$ approximates the true posterior $q_{\text{star-decomp}}(\mathbf{z}_{0}|\mathbf{y})$. The distribution $p_{\bm{\alpha}}$ is decomposed as
$p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})=\sum_{\mathbf{z}_{0}}q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{0})\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y}),$ (6)
@@
where $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ is a categorical distribution parameterized by $\bm{\alpha}$, where for each time step $t$ and dimension $i$, $\bm{\alpha}_{t,i,\cdot}$ is a probability vector in the simplex $\Delta^{K-1}$, defined as $\tilde{p}_{\bm{\alpha}}(z_{0,i}|\mathbf{z}_{t},\mathbf{y})=\text{Cat}\left(z_{0,i};\bm{\alpha}_{t,i,\cdot}\right)$, i.e., $\tilde{p}_{\bm{\alpha}}(z_{0,i}=k|\mathbf{z}_{t},\mathbf{y})=\alpha_{t,i,k}$. In the subsequent discussions and optimization steps, we assume that $\alpha_{t,i,\cdot}$ is always normalized to lie on the simplex. This decomposition stems from the fact that the distribution $q_{\text{star-decomp}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$ ($q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$) can be expressed as $\sum_{\mathbf{z}_{0}}q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{0})q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ based on the conditional independence. Note that both $q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{0})$ and $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ have a mean field structure with independent categorical distributions across dimensions. Consequently, $p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})$, obtained by marginalizing over $\mathbf{z}_{0}$, inherits this mean field property. For notational convenience, we denote the slice of distribution parameter $\bm{\alpha}$ at time step $t$ as $\bm{\alpha}_{t}\in\mathbb{R}^{d_{\mathbf{z}}\times K}$.
@@
To ensure that the marginal distribution of the variational distribution $p_{\bm{\alpha}}$ closely approximates that of $q_{\textnormal{star-decomp}}$, the parameters $\bm{\alpha}$ of the variational distribution are obtained by optimizing an objective function derived from the following theorem:
{restatable}
@@
theoremoptimobject Let $p_{\bm{\alpha}}$ be a distribution with the parameterization given by the decomposition in (6). Then, for any measurements $\mathbf{y}$, the following inequality holds for the KL divergence between the marginal distributions:
@@
$D_{\mathrm{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{y})\|q_{\text{star-decomp}}(\mathbf{z}_{0}|\mathbf{y})\right)\leq\sum_{t=1}^{T}\mathbb{E}_{\mathbf{z}_{t}\sim p_{\bm{\alpha}}(\mathbf{z}_{t}|\mathbf{y})}\left[D_{\mathrm{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)\right],$ (7)
where $p_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{y})$ is the variational marginal distribution parameterized by $\bm{\alpha}$, $q_{\text{star}}(\mathbf{z}_{0}|\mathbf{y})$ is the true posterior distribution, $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ is the variational conditional distribution as defined in (6), $q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ is the true conditional distribution, $p_{\bm{\alpha}}(\mathbf{z}_{t}|\mathbf{y})$ is the marginal distribution at time step $t$, and $T$ is the total number of time steps in the diffusion process.
@@
The proof is provided in the Appendix. Based on this inequality, we aim to minimize each term in the sum on the right-hand side. Since $\tilde{p}_{\bm{\alpha}}$ is a different categorical distribution at each $t$, we minimize $\bm{\alpha}$ for each time step, ultimately aiming to minimize the left-hand side. Each term on the right-hand side of (7) can be decomposed into a term representing deviation from the prior and a term representing likelihood with respect to the observed data:
{restatable}
@@
lemmadecompobj The KL divergence between the variational distribution $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ and the true posterior $q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ can be decomposed into two terms:
@@
$D_{\mathrm{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)=D_{\mathrm{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t})\right)-\mathbb{E}_{\mathbf{z}_{0}\sim\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}\left[\log q_{\text{star}}(\mathbf{y}|\mathbf{z}_{0})\right],$ (8)
@@
This decomposition enables us to separately consider the fit to the prior and the consistency with the measurement data, and we approximate both terms in a computationally tractable form as follows. First, the KL term on the right-hand side of (8) remains intractable. However, we note that the star-shaped noise process shares the conditional distribution $q_{\text{star}}(\mathbf{z}_{t}|\mathbf{z}_{0})$ with the original Markov noise process. Consequently, the reverse conditional distribution $q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t})$ will also be identical for both processes. Since the prior of the pre-trained discrete diffusion models is trained to approximate this distribution, we substitute this prior model $\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})$ for $q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t})$ into the objective function of (8). This substitution transforms the term into a KL divergence between two categorical distributions, enabling the computation of gradients with respect to the parameter $\bm{\alpha}$.
@@
Based on Theorem 3.3 and Lemma 3.3, our proposed inverse problem solving method with discrete diffusion prior, G2D2, optimizes the parameter $\bm{\alpha}$ of $p_{\bm{\alpha}}$ for $t=T,\ldots,1$ while sequentially sampling $\mathbf{z}_{0:T}$. In the optimization step, any continuous optimization method, such as Adam (Kingma, 2014) and RAdam (Liu et al., 2020), can be used. Implementation considerations are discussed in the following section. This algorithm is detailed in Algorithm 1, and G2D2 is illustrated in Figure 1.
@@
5:    Initialize: $\bm{\alpha}_{t}\propto\log\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})$